OpenStudy (anonymous):

Use substitution method y+2y=14 , 2y=x-10

OpenStudy (anonymous):

\[y=\frac{x-10}{2}\]

OpenStudy (anonymous):

plug it into the other

OpenStudy (anonymous):

\[\frac{x-10}{2}+x-10=x-10\]

OpenStudy (anonymous):

I need more help plzI

OpenStudy (anonymous):

everything cancels and x=10 plug ten in for x on the other equation

OpenStudy (anonymous):

So what about y

OpenStudy (anonymous):

if 2y=x-10 and you sub in 10 for where x is what would y equal?

OpenStudy (anonymous):

I have no clue